A quality control specialist for a restaurant chain takes a random sample of size 10 to check the amount of soda served in the 16 oz. serving size. The sample mean is 13.20 with a sample standard deviation of 1.51. Assume the underlying population is normally distributed.
Find the 95% confidence interval for the true population mean for the amount of soda served.
Answer #1

Solution :

t\alpha /2,df = 2.262

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.262 * (1.51 / \sqrt 10)

Margin of error = E = 1.08

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

13.20 - 1.08 < \mu < 13.20 + 1.08

12.12 < \mu < 14.28

(12.12 , 14.28)